A study of working actors looked at age and gender. One sample of 65 male actors had a mean age of 28 and a standard deviation of 3. The other sample included 65 female actors with a mean age of 39 and a standard deviation of 2, Estimate with 98% confidence the difference between the average ages of working male (M1) and female (H2) actors. Round answers to the nearest hundredth.
Answer #1

here as sample size is high we may use z distribution

male female
x1            = 28.000 x2            = 39.000
n1           = 65.000 n2           = 65.000
σ1           = 3.000 σ2           = 2.000
std error of difference σ1-σ2= √(σ21/n122/n2)    = 0.4472
for 98 % CI value of z= 2.3263
margin of error E=z*std error                            = 1.0404
lower confidence bound=estimated mean diff.-margin of error = -12.04
Upper confidence bound=estimated mean diff+margin of error= -9.96

hence 98% CI =-12.04<\mu1-\mu2 <-9.96

(Note from t - distributioon this interval can be (-12.06 ; -9.94) or (-12.05 ; -9.95)